HUBBLE SPACE TELESCOPE DAILY REPORT #1771 PERIOD COVERED: 0000Z (UTC) 12/03/96 - 0000Z (UTC) 12/04/96 Daily Status Report as of 339/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of FOS/BL 6913 (FOS Cycle 6: Location of = Spectra and Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make a = series of stability calibration measurements of the wavelength location = and the location of the spectra. All the observations were internals. = The observations were completed as planned, and no problems were = reported. 1.2 Completed GHRS/1/2 6415 (Thermal Pressures of the Local = Interstellar Medium) The High Resolution Spectrograph (Sides One and Two) was used to = make observations of the star HD186882 in order to study the thermal = pressure of the interstellar medium which lies in the line of sight to = that star. The observations were completed as scheduled, and no = anomalies were reported. 1.3 Completed Two Sets of WF/PC-2 6361 (High Resolution Imaging of = X-Ray Selected AGNs) The WF/PC-2 was used to make high resolution observations of the = known low, red-shifted x-ray active galactic nuclius (AGN) MS1138.1+0. = These data will allow a study of the optical properties of an unbiased = sample of AGN host galaxy. The observations were completed as scheduled, = and no problems were reported. 1.4 Completed FOS/BL 6683 (Eclipsing Binaries in the Magellanic Clouds = - Fundamental Properties A) The Faint Object Spectrograph (Blue Side) was used to make = spectrographic observations of O/B Class, eclipsing binary system star = LMC-HV2274located in the Large Magellanic Cloud. From these data and the = latest Kurucz model atmospheres, accurate temperatures will be derived. = The observations were executed as scheduled, and no anomalies were noted. = 1.5 Completed COSTAR/U 6895 (FOC Cycle 6 PSF Monitoring) The Faint Object Camera/COSTAR were used to make observations of = object 01-001 (aka, white dwarf star BPM16274) in order to check out the = point spread function (PSF) and focus of the system. An internal flat = calibration image was also taken. The observations were executed as = scheduled, and no problems were reported. 1.6 Completed Two Sets of WF/PC-2 6802 (Cosmology With the Medium Deep = Survey) The WF/PC-2 was used to make observations taken in parallel with = the prime science observations as part of the Medium Deep Sky Survey. = The observations were executed as scheduled, and no problems were = reported. 1.7 Completed GHRS/2 6599 (Boron in Halo Giants: A Solution to the D+3 = He Crisis in Big Bang Nucleus) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the boron abundances of giant halo star = HD63791. Comparison of the abundances among other stars of the same type = will allow the investigator to determine whether mass loss or mixing = causes the low Carbon-12/ Carbon-13 ratios seen in these type stars. The = observations were executed as scheduled, and no problems were reported. 1.8 Completed WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were reporte= d. 1.9 Completed FOS/BL 6532 (A Definitive Test for the Excitation = Mechanism of LINERs) The Faint Object Spectrograph (Blue Side) was used to make = ultraviolet and visible spectrographic observations of the spiral galaxy = NGC-3031 (aka, M-81) as part of a definitive test for the external = excitation mechanism of LINERs. This is to determine if LINERs are = either photoionized or shock ionized, by obtaining FOS spectra from 1150 = A0 through 6800 A0 by viewing its gas disk. The observations were execute= d as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science observati= ons. The intent is to search for extra-galactic sources with flux = excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The = observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: Aside from several temperature limit violations which regularly = occur at this time of the year, there were no significant problems = reported. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH